# Hashtag Re-Appropriation for Audience Control on Recommendation-Driven Social Media Xiaohongshu (rednote)

### v1

@@ Abstract.
Algorithms have played a central role in personalized recommendations on social media. However, they also present significant obstacles for content creators trying to predict and manage their audience reach. This issue is particularly challenging for marginalized groups seeking to maintain safe spaces. Our study explores how women on Xiaohongshu (rednote), a recommendation-driven social platform, proactively re-appropriate hashtags (e.g., #{CJK}UTF8gbsn宝宝辅食, Baby Supplemental Food) by using them in posts unrelated to their literal meaning. The hashtags were strategically chosen from topics that would be uninteresting to the male audience they wanted to block. Through a mixed-methods approach, we analyzed the practice of hashtag re-appropriation based on 5,800 collected posts and interviewed 24 active users from diverse backgrounds to uncover users’ motivations and reactions towards the re-appropriation. This practice highlights how users can reclaim agency over content distribution on recommendation-driven platforms, offering insights into self-governance within algorithmic-centered power structures.
@@
journalyear: 2025copyright: rightsretainedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713379isbn: 979-8-4007-1394-1/25/04submissionid: 6140ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Moreover, the opaque nature of recommendation mechanisms presents unique challenges for groups such as women, Black people, and LGBTQ+ individuals (Klassen et al., 2021; Karizat et al., 2021). Karizat et al. (2021) found that TikTok’s “For You” Page algorithm actively suppresses content related to marginalized social identities. Experienced the restriction of content visibility on TikTok, content creators increasingly invent words to bypass content moderation algorithms from banning their videos, such as using ‘le\$bian’ for ‘lesbian’  (Steen et al., 2023). Also, the lack of direct control over content visibility could result in “context collapse”, when content intended for a single audience is exposed to diverse audiences within one environment, leading to unexpected or even negative consequences for their self-presentation (Davis and Jurgenson, 2014; Marwick and Boyd, 2011). For example, on Black Twitter, context collapse occurs when tweets are accidentally exposed to audiences outside of the Black community, making it difficult for users to manage self-presentation and avoid miscommunication  (Klassen et al., 2021; Allen, 2023). Given algorithms’ opaqueness, users have limited ways to predict or control their audience (DeVito et al., 2018; DeVito, 2021; DeVito et al., 2017b, a). Content creators can add hashtags to extend their content, while content consumers can use hashtags as keywords to search content; recommendation algorithms leverage hashtags to deliver posts to targeted audiences  (DeVito et al., 2018; TikTok Support, [n. d.]), further shaping the visibility dynamics. Recognizing the capability of hashtags, users have developed increasingly creative approaches to optimize their use.
@@
In this paper, we investigated the widespread use of the hashtag #{CJK}UTF8gbsn宝宝辅食 (Baby Supplemental Food, abbreviated as BSF) [^1] The quote shows the original Chinese hashtags and its English translation and abbreviation in the parentheses. on the Chinese social media Xiaohongshu [^2] Xiaohongshu means “Little Red Book,” also known as rednote in English. It is a social sharing and e-commerce platform. As of 2024, Xiaohongshu has over 209 million monthly active users  (Statista, 2024b). https://www.xiaohongshu.com/. #BSF became viral in the Xiaohongshu community for its unexpected and diverse usage, which has been widely reported and discussed online  (Sohu, 2024; The Paper (2024), Pengpai News; Xiaohongshu, 2024)  [^3] These articles are: {CJK}UTF8gbsn宝宝辅食与小红书隔“男”墙 (Baby Supplemental Food and Xiaohongshu’s wall against “men”) on Sohu, “{CJK}UTF8gbsn宝宝辅食”tag:互联网新型护盾 (“Baby Supplemental Food” tag: a new type of internet shield) on Pengpai, and {CJK}UTF8gbsn新传热点宝宝辅食?更适合躲避男凝的标签 (News communication hot topic Baby Supplemental Food? a better tag to avoid the male gaze) on Xiaohongshu.. The articles reported many Xiaohongshu female users re-appropriate #BSF in an attempt to moderate the recommendation algorithm not push their posts to male users, instead of tagging baby food-related posts. The hashtag re-appropriation practice reflects women’s effort to manage their safe spaces under the challenges of recommendation-driven platforms like Xiaohongshu. Driven by this initial observation, we systematically investigate the phenomenon from a mixed-methods approach. Unlike network-based platforms such as Facebook and Instagram, where content delivery primarily depends on user social networks, Xiaohongshu delivers content across the entire platform based on user behavior and interests (Zhang et al., 2024), making audiences beyond their direct followers or group memberships (Cao et al., 2024; Johnson et al., 2022). On the other hand, in contrast to community-based platforms like Reddit—where user engagement is governed by community rules and moderators, behaviors on Xiaohongshu are regulated mainly by general platform rules and moderation (Xiaohongshu, 2023). These differences raise the question of how users can gain more control over algorithms—both in terms of what they see and who sees their posts—and maintain safe spaces on recommendation-driven platforms. Building on prior literature on hashtag activism (Mueller et al., 2021; Yang, 2016), feminism (D’ignazio and Klein, 2023; Bardzell, 2010), and user control in algorithmic platforms (Witzenberger, 2018; Van der Nagel, 2018; Burrell et al., 2019a), our study aims to address the following research questions:
@@
Our research reveals the practice of #BSF hashtag re-appropriation as a process of blocking, attracting, and evolving. Due to ineffective reporting and blocking mechanisms, and a loss of trust in the platform’s women-user-friendly stance, women began to adopt hashtag re-appropriation as a strategy. Different user groups responded to this re-appropriation with varying degrees of support and misunderstanding. The case of hashtag re-appropriation enabled women to shape their narratives through interconnected posts while simultaneously blocking unwanted viewers, demonstrating a form of everyday resistance to digital feminism in China. It also inspires the broader implications of audience control on recommendation-driven platforms and the potential for user self-governance within algorithm-centered structures.
@@ 2.1. Digital Feminism Practices in China
As digital feminism develops, anti-feminists have also become active on Chinese social media. They use the derogatory term “{CJK}UTF8gbsn田园女权 (Rural Feminism)” to attack women users (Yang et al., 2023; Wu and Dong, 2019), creating a false dichotomy between “rural” and “authentic” feminism. They condemn certain feminist demands (such as men taking on more household responsibilities) while endorsing “moderate” claims (such as women rejecting bride prices) (Yang et al., 2023). Opposition to feminism also exists among women, who criticize it as a hidden disciplinary force that undermines female solidarity (such as wearing makeup) (Wang and Chang, 2023). Some also view digital feminism as merely a trendy business (Wang and Chang, 2023). Additionally, a large nuanced subgroup exists distinct from feminists or anti-feminists.  Yin and Sun (2021) used the term “pro-change counterpublics” to describe those who refuse to identify as feminists but actively protest against sexual harassment in online discussions. This subgroup differs from those who fully embrace feminist ideas.
@@ 4.2.1. Data Collection
We collected posts with the original #BSF hashtag and its derived hashtags from Xiaohongshu from February to April 2024 through web scraping. For each post, we collected its title, content, and hashtags. As #BSF gained popularity, its usage became more complex. Additional hashtags with similar purposes, such as {CJK}UTF8bsmi#寶寶輔食 (BSF in Traditional Chinese, abbreviated as BSF(TC)) and {CJK}UTF8gbsn#男性结扎 (Male Sterilization, abbreviated as MS), also emerged  (Sohu, 2024). Then, we identified three popular hashtags derived from #BSF through empirical observations of user discussions and co-use with #BSF on Xiaohongshu, and collected posts using these hashtags. The derived hashtags include #BSF(TC), #MS, and {CJK}UTF8gbsn#185大帅哥 (185 Handsome Guy, abbreviated as HG).
We also constructed a baseline dataset that represents the general post content on Xiaohongshu. We selected three hashtags with high exposures and active users—{CJK}UTF8gbsn#美食 (Cuisine), {CJK}UTF8gbsn#我的日常 (My Daily), and {CJK}UTF8gbsn#不懂就问有问必答 (Ask if you don’t know, answer if someone asks)—and collected their posts. These three hashtags correspond to three popular types of posts we observed on Xiaohongshu: interest, daily sharing, and Q&A. These posts were then compared with those containing #BSF and its derived hashtags in the following analysis. Table 1 provides a summary of all selected hashtags and collected posts.
@@
Original {CJK}UTF8gbsn宝宝辅食 Baby Supplemental Food, BSF 29.75 3,311
Derived {CJK}UTF8bsmi寶寶輔食 BSF in Traditional Chinese, BSF(TC) 0.083 324
{CJK}UTF8gbsn185大帅哥 185 Handsome Guy, HG 2.88 613
{CJK}UTF8gbsn男性结扎 Male Sterilization, MS 0.2 338
Baseline {CJK}UTF8gbsn美食 Cuisine 58.04 450
{CJK}UTF8gbsn我的日常 My Daily 81.27 455
{CJK}UTF8gbsn不懂就问有问必答 Ask if you don’t know, answer if someone asks 37.38 309
@@ 4.3. Qualitative Interview Study with Xiaohongshu Users
We conducted 24 semi-structured online interviews from June to July 2024 to understand Xiaohongshu users’ perspectives on the usage of the #BSF and its derived hashtags. Our study was approved by our institution’s Institutional Review Board (IRB).
@@ 4.3.1. Recruitment
To include comprehensive perspectives, we recruited diverse participants, including 17 female (70.83%) and seven male users (29.17%) following the users’ gender ratio on Xiaohongshu. Among them, 4 participants have children; 5 participants disclosed their LGBTQ+ identity; 10 participants have experience in either operating marketing content for companies or operating personal content as self-identified independent content creators. More details about the participants are presented in Table 2.
@@ 5. Findings
Given the mixed-method analysis, our findings reveal hashtag re-appropriation’s evolution process, underlying motivations, diverse user groups’ reactions, also the comparative affordance of hashtags in audience management and content relevance.
@@ 5.1.3. Evolution of Hashtag Re-Appropriation
Further, our co-occurrence network analysis for #BSF and its derived hashtags depicts a more granular usage pattern of derived hashtags. As shown in Figure 5, #BSF, positioned at the core of this network, serves as a primary node, closely associated with each derived hashtag. Each derived hashtag has distinct co-occurring hashtag clusters. For instance, #HG has a co-occurring hashtag cluster about “{CJK}UTF8gbsn油腻文学 (Greasy Literature)”, a trendy term in the online community used by people (mostly women) to make fun of men’s greasy flirts. #MS has two main co-occurring hashtag clusters: one focuses on “{CJK}UTF8gbsn乙女游戏 (Otome Game)”, a genre of female-targeted romance games, and the other on “{CJK}UTF8gbsn妈妈辅食 (Mom Supplemental Food)”. Otome Game players enjoy engaging in romantic experiences with virtual male characters within games. They use #Otome Game to share their game experiences with other peer women players on Xiaohongshu while using #MS to avoid unwanted attention from male users who usually stigmatize women players. #Mom Supplemental Food is also a re-appropriated hashtag, echoing #BSF. Notably, posts using #MS and #Mom Supplemental Food often feature thirst-trap-like photos. Additionally, #BSF(TC) shows a closer connection with #MS than with #HG. #Le (Lesbian) connects #BSF(TC) with #MS, while #Notes Inspiration connects #BSF(TC) with #HG. The hashtag co-occurrence network illustrates how various user groups favor different derived hashtags to achieve their audience control goals and align with their nuanced interests.
@@ 5.2. Motivations of Hashtag Re-Appropriation
Analyzing the usage of #BSF and its derived hashtags paved the way for understanding the underlying reasons behind users’ participation in this re-appropriation. Our interviews revealed three primary challenges users faced on Xiaohongshu that led to hashtag re-appropriation: 1) a shifting user base on the platform, 2) persistent gender-based harassment, and 3) the ineffectiveness of existing safety features.
@@ 5.4.1. Regular Usage
The most common usage is “Attract & Relevant”, that is, using hashtags relevant to the post to attract audiences who are interested in the topic. For example, a user can share a homemade dinner using #{CJK}UTF8gbsn美食 (cuisine) so that the post will be recommended to users who are interested in dinner recipes. Users looking for cuisine in a particular city can search for this hashtag along with the name of the city to find reviews of local restaurants. Specific hashtags relevant to the posts can also be used to attract the audience. For example, #Otome Game, as mentioned before, is used by players with the same interest to find each other in the hashtag topic board. The original use of #BSF also falls into this category, attracting parents of infants who wanted to learn how to make baby supplemental food (Section 5.1.1).

### v2

@@ Abstract.
Algorithms have played a central role in personalized recommendations on social media. However, they also present significant obstacles for content creators trying to predict and manage their audience reach. This issue is particularly challenging for marginalized groups seeking to maintain safe spaces. Our study explores how women on Xiaohongshu (rednote), a recommendation-driven social platform, proactively re-appropriate hashtags (e.g., #宝宝辅食, Baby Supplemental Food) by using them in posts unrelated to their literal meaning. The hashtags were strategically chosen from topics that would be uninteresting to the male audience they wanted to block. Through a mixed-methods approach, we analyzed the practice of hashtag re-appropriation based on 5,800 collected posts and interviewed 24 active users from diverse backgrounds to uncover users’ motivations and reactions towards the re-appropriation. This practice highlights how users can reclaim agency over content distribution on recommendation-driven platforms, offering insights into self-governance within algorithmic-centered power structures.
@@
journalyear: 2025copyright: rightsretainedconference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713379isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing Empirical studies in collaborative and social computing
@@ 1. Introduction
Moreover, the opaque nature of recommendation mechanisms presents unique challenges for groups such as women, Black people, and LGBTQ+ individuals (Klassen et al., 2021; Karizat et al., 2021). Karizat et al. (2021) found that TikTok’s “For You” Page algorithm actively suppresses content related to marginalized social identities. Experienced the restriction of content visibility on TikTok, content creators increasingly invent words to bypass content moderation algorithms from banning their videos, such as using “le\$bian” for “lesbian”  (Steen et al., 2023). Also, the lack of direct control over content visibility could result in “context collapse,” when content intended for a single audience is exposed to diverse audiences within one environment, leading to unexpected or even negative consequences for their self-presentation (Davis and Jurgenson, 2014; Marwick and Boyd, 2011). For example, on Black Twitter, context collapse occurs when tweets are accidentally exposed to audiences outside of the Black community, making it difficult for users to manage self-presentation and avoid miscommunication  (Klassen et al., 2021; Allen, 2023). Given algorithms’ opaqueness, users have limited ways to predict or control their audience (DeVito et al., 2018; DeVito, 2021; DeVito et al., 2017b, a). Content creators can add hashtags to extend their content, while content consumers can use hashtags as keywords to search content; recommendation algorithms leverage hashtags to deliver posts to targeted audiences  (DeVito et al., 2018; TikTok Support, [n. d.]), further shaping the visibility dynamics. Recognizing the capability of hashtags, users have developed increasingly creative approaches to optimize their use.
@@
In this paper, we investigated the widespread use of the hashtag #{CJK*}UTF8gbsn宝宝辅食 (Baby Supplemental Food, abbreviated as BSF) [^1] The quote shows the original Chinese hashtags and its English translation and abbreviation in the parentheses. on the Chinese social media Xiaohongshu [^2] Xiaohongshu means “Little Red Book,” also known as rednote in English. It is a social sharing and e-commerce platform. As of 2024, Xiaohongshu has over 209 million monthly active users  (Statista, 2024b). https://www.xiaohongshu.com/. #BSF became viral in the Xiaohongshu community for its unexpected and diverse usage, which has been widely reported and discussed online  (Sohu, 2024; The Paper (2024), Pengpai News; Xiaohongshu, 2024)  [^3] These articles are: {CJK*}UTF8gbsn宝宝辅食与小红书隔“男”墙 (Baby Supplemental Food and Xiaohongshu’s wall against “men”) on Sohu, “{CJK*}UTF8gbsn宝宝辅食”tag:互联网新型护盾 (“Baby Supplemental Food” tag: a new type of internet shield) on Pengpai, and {CJK*}UTF8gbsn新传热点宝宝辅食?更适合躲避男凝的标签 (News communication hot topic Baby Supplemental Food? a better tag to avoid the male gaze) on Xiaohongshu.. The articles reported many Xiaohongshu female users re-appropriate #BSF in an attempt to moderate the recommendation algorithm not push their posts to male users, instead of tagging baby food-related posts. The hashtag re-appropriation practice reflects women’s effort to manage their safe spaces under the challenges of recommendation-driven platforms like Xiaohongshu. Driven by this initial observation, we systematically investigate the phenomenon from a mixed-methods approach. Unlike network-based platforms such as Facebook and Instagram, where content delivery primarily depends on user social networks, Xiaohongshu delivers content across the entire platform based on user behavior and interests (Zhang et al., 2024), making audiences beyond their direct followers or group memberships (Cao et al., 2024; Johnson et al., 2022). On the other hand, in contrast to community-based platforms like Reddit—where user engagement is governed by community rules and moderators, behaviors on Xiaohongshu are regulated mainly by general platform rules and moderation (Xiaohongshu, 2023). These differences raise the question of how users can gain more control over algorithms—both in terms of what they see and who sees their posts—and maintain safe spaces on recommendation-driven platforms. Building on prior literature on hashtag activism (Mueller et al., 2021; Yang, 2016), feminism (D’ignazio and Klein, 2023; Bardzell, 2010), and user control in algorithmic platforms (Witzenberger, 2018; Van der Nagel, 2018; Burrell et al., 2019a), our study aims to address the following research questions:
@@
Our research reveals the practice of #BSF hashtag re-appropriation as a process of blocking, attracting, and evolving. Due to ineffective reporting and blocking mechanisms and a loss of trust in the platform’s women-user-friendly stance, women began to adopt hashtag re-appropriation as a strategy. Different user groups responded to this re-appropriation with varying degrees of support and misunderstanding. The case of hashtag re-appropriation enabled women to shape their narratives through interconnected posts while simultaneously blocking unwanted viewers, demonstrating a form of everyday resistance to digital feminism in China. It also inspires the broader implications of audience control on recommendation-driven platforms and the potential for user self-governance within algorithm-centered structures.
@@ 2.1. Digital Feminism Practices in China
As digital feminism develops, anti-feminists have also become active on Chinese social media. They use the derogatory term “{CJK*}UTF8gbsn田园女权 (Rural Feminism)” to attack women users (Yang et al., 2023; Wu and Dong, 2019), creating a false dichotomy between “rural” and “authentic” feminism. They condemn certain feminist demands (such as men taking on more household responsibilities) while endorsing “moderate” claims (such as women rejecting bride prices) (Yang et al., 2023). Opposition to feminism also exists among women, who criticize it as a hidden disciplinary force that undermines female solidarity (such as wearing makeup) (Wang and Chang, 2023). Some also view digital feminism as merely a trendy business (Wang and Chang, 2023). Additionally, a large nuanced subgroup exists distinct from feminists or anti-feminists.  Yin and Sun (2021) used the term “pro-change counterpublics” to describe those who refuse to identify as feminists but actively protest against sexual harassment in online discussions. This subgroup differs from those who fully embrace feminist ideas.
@@ 4.2.1. Data Collection
We collected posts with the original #BSF hashtag and its derived hashtags from Xiaohongshu from February to April 2024 through web scraping. For each post, we collected its title, content, and hashtags. As #BSF gained popularity, its usage became more complex. Additional hashtags with similar purposes, such as {CJK*}UTF8bsmi#寶寶輔食 (BSF in Traditional Chinese, abbreviated as BSF(TC)) and {CJK*}UTF8gbsn#男性结扎 (Male Sterilization, abbreviated as MS), also emerged  (Sohu, 2024). Then, we identified three popular hashtags derived from #BSF through empirical observations of user discussions and co-use with #BSF on Xiaohongshu, and collected posts using these hashtags. The derived hashtags include #BSF(TC), #MS, and {CJK*}UTF8gbsn#185大帅哥 (185 Handsome Guy, abbreviated as HG).
We also constructed a baseline dataset that represents the general post content on Xiaohongshu. We selected three hashtags with high exposures and active users—{CJK*}UTF8gbsn#美食 (Cuisine), {CJK*}UTF8gbsn#我的日常 (My Daily), and {CJK*}UTF8gbsn#不懂就问有问必答 (Ask if you don’t know, answer if someone asks)—and collected their posts. These three hashtags correspond to three popular types of posts we observed on Xiaohongshu: interest, daily sharing, and Q&A. These posts were then compared with those containing #BSF and its derived hashtags in the following analysis. Table 1 provides a summary of all selected hashtags and collected posts.
@@
Original {CJK*}UTF8gbsn宝宝辅食 Baby Supplemental Food, BSF 29.75 3,311
Derived {CJK*}UTF8bsmi寶寶輔食 BSF in Traditional Chinese, BSF(TC) 0.083 324
{CJK*}UTF8gbsn185大帅哥 185 Handsome Guy, HG 2.88 613
{CJK*}UTF8gbsn男性结扎 Male Sterilization, MS 0.2 338
Baseline {CJK*}UTF8gbsn美食 Cuisine 58.04 450
{CJK*}UTF8gbsn我的日常 My Daily 81.27 455
{CJK*}UTF8gbsn不懂就问有问必答 Ask if you don’t know, answer if someone asks 37.38 309
@@ 4.3. Qualitative Interview Study with Xiaohongshu Users
We conducted 24 semi-structured online interviews from June to July 2024 to understand Xiaohongshu users’ perspectives on the usage of the #BSF and its derived hashtags. Our study was approved by our institution’s Institutional Review Board.
@@ 4.3.1. Recruitment
To include comprehensive perspectives, we recruited diverse participants, including 17 female (70.83%) and seven male users (29.17%) following the users’ gender ratio on Xiaohongshu. Among them, 4 participants have children; 5 participants disclosed their LGBTQ+ identity; 10 participants have experience in either operating marketing content for companies or operating personal content as self-identified independent content creators. We refer to participants’ sexual orientations based on their self-identification, including only those who voluntarily disclosed this information during interviews. Additionally, we obtained their explicit consent for its inclusion in the paper. More details about the participants are presented in Table 2.
@@ 5. Findings
Given the mixed-method analysis, our findings reveal hashtag re-appropriation’s evolution process, underlying motivations, diverse user groups’ reactions, as well as the comparative affordance of hashtags in audience management and content relevance.
@@ 5.1.3. Evolution of Hashtag Re-Appropriation
Further, our co-occurrence network analysis for #BSF and its derived hashtags depicts a more granular usage pattern of derived hashtags. As shown in Figure 5, #BSF, positioned at the core of this network, serves as a primary node, closely associated with each derived hashtag. Each derived hashtag has distinct co-occurring hashtag clusters. For instance, #HG has a co-occurring hashtag cluster about “{CJK*}UTF8gbsn油腻文学 (Greasy Literature)”, a trendy term in the online community used by people (mostly women) to make fun of men’s greasy flirts. #MS has two main co-occurring hashtag clusters: one focuses on “{CJK*}UTF8gbsn乙女游戏 (Otome Game)”, a genre of female-targeted romance games, and the other on “{CJK*}UTF8gbsn妈妈辅食 (Mom Supplemental Food)”. Otome Game players enjoy engaging in romantic experiences with virtual male characters within games. They use #Otome Game to share their game experiences with other peer women players on Xiaohongshu while using #MS to avoid unwanted attention from male users who usually stigmatize women players. #Mom Supplemental Food is also a re-appropriated hashtag, echoing #BSF. Notably, posts using #MS and #Mom Supplemental Food often feature thirst-trap-like photos. Additionally, #BSF(TC) shows a closer connection with #MS than with #HG. #Le (Lesbian) connects #BSF(TC) with #MS, while #Notes Inspiration connects #BSF(TC) with #HG. The hashtag co-occurrence network illustrates how various user groups favor different derived hashtags to achieve their audience control goals and align with their nuanced interests.
@@ 5.2. Motivations of Hashtag Re-Appropriation
Analyzing the usage of #BSF and its derived hashtags paved the way for understanding the underlying reasons behind users’ participation in this re-appropriation. Our interviews reveal three primary challenges users faced on Xiaohongshu that led to hashtag re-appropriation: 1) a shifting user base on the platform, 2) persistent gender-based harassment, and 3) the ineffectiveness of existing safety features.
@@ 5.4.1. Regular Usage
The most common usage is “Attract & Relevant”, that is, using hashtags relevant to the post to attract audiences who are interested in the topic. For example, a user can share a homemade dinner using #{CJK*}UTF8gbsn美食 (cuisine) so that the post will be recommended to users who are interested in dinner recipes. Users looking for cuisine in a particular city can search for this hashtag along with the name of the city to find reviews of local restaurants. Specific hashtags relevant to the posts can also be used to attract the audience. For example, #Otome Game, as mentioned before, is used by players with the same interest to find each other in the hashtag topic board. The original use of #BSF also falls into this category, attracting parents of infants who wanted to learn how to make baby supplemental food (Section 5.1.1).